OpenStudy (hanner_b_nanner):

Help please? Fan&&Medal Is rate of change, and changing the constant the same thing? I'm confused...

OpenStudy (amistre64):

is the a more full question that this is from?

OpenStudy (amistre64):

rate of change has to do with slopes, constants have to do with vertical displacements

OpenStudy (hanner_b_nanner):

I have a function f(x)=2(x-11)^2 +3 , and my question is Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (amistre64):

oh, then you want to find the average rate of change (the slope of the line) between the stated points.

OpenStudy (amistre64):

find (4,f(4)) and (7,f(7)) then determine the slope of the line that connects them

OpenStudy (aum):

Average rate of change for f(x) from x = 4 to x = 7 = \(\Large \frac{f(7) - f(4)}{7-4} = ?\)

OpenStudy (hanner_b_nanner):

Sooo, I plug in (4,f(4)) and (7,f(7)) to my function?? Or..

OpenStudy (amistre64):

f(a) = 2(a-11)^2 + 3 f(b) = 2(b-11)^2 + 3 f(a) - f(b) [2(a-11)^2 + 3] - [2(b-11)^2 + 3] 2(a-11)^2 + 3 -2(b-11)^2 - 3 2(a-11)^2 -2(b-11)^2 2 [(a-11)^2 - (b-11)^2] now we have a difference of squares ... 2 [(a-11) - (b-11)] [(a-11) + (b-11)] 2 (a-11-b+11) (a-11+b-11) 2 (a-b) (a+b)

OpenStudy (amistre64):

well, yeah. using the slope formula, or definition the change in f(x) as x moves from 4 to 7, divided by the change from 4 to 7 \[\frac{f(4)-f(7)}{4-7}\]

OpenStudy (amistre64):

ironically, we have a-b on the denominator so this simplifies to just 2(a+b)

OpenStudy (hanner_b_nanner):

ohhh, okay!! :)) Thank you!